# Ringing out General Relativity: Quasi-normal mode frequencies for black holes of any spin in modified gravity

### v1

@@ Abstract
For future parameter estimation of detected ringdown signals, we construct the optimal fitting polynomial and find that it is of degree seven and has coefficients $w_{j=0,\ldots,7}$, given in Table 1. The $w_{0},w_{1}$ and $w_{2}$ coefficients for the polar mode are consistent with those presented in Pierini and Gualtieri (2022), despite their use of a second-order small-spin expansion, and different numerical and fitting algorithms. Using this fitting function, we estimate that, if a ringdown signal emitted by a remnant black hole with $a=0.7$ in general relativity is detected by LIGO-Virgo-KAGRA with a relative measurement uncertainty of $\delta\hat{f}_{022}\sim 0.05$ Abbott et al. (2021), the METRICS frequencies computed here should allow the conservative constraint $\zeta\lesssim\delta\hat{f}_{022}/\min(|\text{Re}~{}\omega^{(1)}_{\rm A}|,|\text{Re}~{}\omega^{(1)}_{\rm P}|)\sim{\cal{O}}(0.5)$, which translates to a constraint of $\sqrt{\alpha}\lesssim 12$ km for a black hole with mass $10M_{\odot}$.
@@
0 $0.0598144+0.00591626i~{}(0.0008+0.0003i)$ $-0.224292-0.0754365i~{}(0.0003+0.00006i)$
1 $-0.00755503++0.27555i~{}(0.03+0.07i)$ $-0.481923+0.0400771i~{}(0.03+0.01i)$
2 $2.83089-4.20805i~{}(0.5+1.3i)$ $2.72878-0.497164i~{}(0.7+0.2i)$
3 $-21.7428+32.7515i~{}(3.1+9.9i)$ $-25.1868+6.96795i~{}(5.2+1.6i)$
4 $89.1059-129.867i~{}(9.9+35.5i)$ $99.3176-24.6157i~{}(19.2+5.2i)$
5 $-191.851++264.05i~{}(16.8+65.4i)$ $-202.491+48.2059i~{}(36.2+8.9i)$
6 $204.863-268.745i~{}(14.3+59.8i)$ $208.182-47.6168i~{}(33.7+7.5i)$
7 $-86.7247+108.37i~{}(4.82+21.5i)$ $-85.7465+17.8104i~{}(12.3+2.5i)$
Table 1: Numerical values and uncertainties (in parentheses, rounded to the nearest significant decimal) of the coefficients of the optimal fitting polynomial of the axial and polar frequencies, using the fitting function $M\omega^{(1)}=\sum_{j=0}^{7}w_{j}a^{j}$.
@@
The extended METRICS framework can be applied to other theories of gravity and other scenarios. We have here focused on theories with massless scalar fields, where the background metric and scalar are time-independent, and which only have two (transverse-traceless) gravitational wave polarizations Wagle et al. (2019). Nonetheless, the extended METRICS approach should be able to handle massive fields (with the appropriate modification to $V(\Phi)$ and the boundary conditions of the fields), as well as dynamical vector fields and scalar or vector polarizations, through the appropriate generalization of the field equations and of the initial choices of METRICS iterations respectively. METRICS can also be applied to calculate the complex frequencies of Kerr black holes immersed in a dark matter or accretion disk environment. The environmental modifications to the black hole background and any dynamical environmental interactions encoded in the field equations can be treated within the extended METRICS approach. These are just but a few examples of the pathways that METRICS creates for future work.
@@ Acknowledgements.
The authors acknowledge the support from the Simons Foundation through Award No. 896696, the NSF through award PHY-2207650 and NASA through Grant No. 80NSSC22K0806. A.K.W.C would like to thank Emanuele Berti and Dongjun Li for insightful discussion. N.Y. would like to thank Takahiro Tanaka for insightful discussion about this work. The authors would like to thank Alessandra Buonanno, Vitor Cardoso, Gregorio Carullo, Gregory Gabadadze, Leonardo Gualtieri and Shinji Tsujikawa for comments on the initial version of the manuscript. The calculations and results reported in this paper were produced using the computational resources of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with NCSA, and is supported by funds from the University of Illinois at Urbana-Champaign. The author would like to specially thank the investors of the IlliniComputes initiatives and GravityTheory computational nodes for permitting the authors to execute runs related to this work using the relevant computational resources.

### v3

@@ Abstract
For future parameter estimation of detected ringdown signals, we construct the optimal fitting polynomial and find that it is of degree seven and has coefficients $w_{j=0,\ldots,8}$, given in Table 1. The $w_{0},w_{1}$, and $w_{2}$ coefficients for the polar mode are consistent with those presented in Pierini and Gualtieri (2022), despite their use of a second-order small-spin expansion, and different numerical and fitting algorithms. Using this fitting function, we estimate that, if a ringdown signal emitted by a remnant black hole with $a=0.7$ in general relativity is detected by LIGO-Virgo-KAGRA with a relative measurement uncertainty of $\delta\hat{f}_{022}\sim 0.05$ Abbott et al. (2021), the METRICS frequencies computed here should allow the conservative constraint $\zeta\lesssim\delta\hat{f}_{022}/\min(|\text{Re}~{}\omega^{(1)}_{\rm A}|,|\text{Re}~{}\omega^{(1)}_{\rm P}|)\sim{\cal{O}}(0.5)$, which translates to a constraint of $\sqrt{\alpha}\lesssim 12$ km for a black hole with mass $10M_{\odot}$.
@@
0 $0.055241+0.00686713i~{}(9.53\times 10^{-6}+0.00686713i)$ $-0.215202-0.0734094i~{}(0.00457066+0.00233301i)$
1 $0.985294+0.0636233i~{}(0.891167+0.0013926i)$ $-2.51816-0.411031i~{}(0.892279+0.271754i)$
2 $-18.4902+0.325569i~{}(18.1278+0.029502i)$ $48.4659+9.59898i~{}(19.3688+5.6713i)$
3 $157.802-4.46276i~{}(146.535+0.250497i)$ $-431.428-82.4765i~{}(163.664+47.2939i)$
4 $-682.224+19.8274i~{}(609.58+1.09047i)$ $1935.01+378.832i~{}(692.595+202.145i)$
5 $1660.08-55.4163i~{}(1429.63+2.65319i)$ $-4831.15-965.058i~{}(1614.95+484.046i)$
6 $-2307.78+87.744i~{}(1910.19+3.64319i)$ $6808.99+1386.33i~{}(2108.76+656.174i)$
7 $1711.32-73.0696i~{}(1356.36+2.63502i)$ $-5065.1-1050.91i~{}(1445.54+470.048i)$
8 $-526.516+25.0324i~{}(396.962+0.779457i)$ $1544.32+325.905i~{}(405.09+138.139i)$
Table 1: Numerical values and uncertainties (in parentheses, rounded to the nearest significant decimal) of the coefficients of the optimal fitting polynomial of the axial and polar frequencies, using the fitting function $M\omega^{(1)}=\sum_{j=0}^{8}w_{j}a^{j}$.
@@
The extended METRICS framework can be applied to other theories of gravity and other scenarios. We have here focused on theories with massless scalar fields, where the background metric and scalar are time independent, and which only have two (transverse-traceless) gravitational wave polarizations Wagle et al. (2019). Nonetheless, the extended METRICS approach should be able to handle massive fields (with the appropriate modification to $V(\Phi)$ and the boundary conditions of the fields), as well as dynamical vector fields and scalar or vector polarizations, through the appropriate generalization of the field equations and of the initial choices of METRICS iterations, respectively. METRICS can also be applied to calculate the complex frequencies of Kerr black holes immersed in a dark matter or accretion disk environment. The environmental modifications to the black hole background and any dynamical environmental interactions encoded in the field equations can be treated within the extended METRICS approach. These are just but a few examples of the pathways that METRICS creates for future work.
@@ Acknowledgements.
The authors acknowledge the support from the Simons Foundation through Award No. 896696, the NSF through Grant No. PHY-2207650 and NASA through Grant No. 80NSSC22K0806. A. K. W. C would like to thank Emanuele Berti and Dongjun Li for insightful discussion. N.Y. would like to thank Takahiro Tanaka for insightful discussion about this work. The authors would like to thank Alessandra Buonanno, Vitor Cardoso, Gregorio Carullo, Gregory Gabadadze, Leonardo Gualtieri and Shinji Tsujikawa for comments on the initial version of the manuscript. The calculations and results reported in this Letter were produced using the computational resources of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with National Center for Supercomputing Applications (NCSA), and is supported by funds from the University of Illinois at Urbana-Champaign, and used Delta at NCSA through allocation PHY240142 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services &\&& Support (ACCESS) program, which is supported by National Science Foundation Grants No. 2138259, No. 2138286, No. 2138307, No. 2137603, and No. 2138296. The author would like to specially thank the investors of the IlliniComputes initiatives and GravityTheory computational nodes for permitting the authors to execute runs related to this work using the relevant computational resources.